# FactReasoner: A Probabilistic Approach to Long-Form Factuality Assessment for Large Language Models

### v1

@@ 3.3 The FactReasoner Pipeline and Variants
The proposed FactReasoner pipeline for long-form factuality assessment is shown in Figure 3 and consists of four main stages called Atomizer, Reviser, Retriever and Evaluator, respectively. It takes as input a response $y$ and outputs the marginal posterior probabilities $P(a_{i})$ of $y$’s atomic units together with the factuality measures described earlier, such as $Pr(y)$, $F_{1}@K(y)$ and $\mathcal{E}(y)$, respectively.
@@
FactReasoner 1 (FR1).
In this case, for each atom variable $A_{i}$ up to $k$ most relevant contexts $\{C_{1}^{i},...,C_{k}^{i}\}$ are retrieved and only the relationships between atom $A_{i}$ and its corresponding contexts are considered, namely only the factors $f(A_{i},C_{j}^{i})$ are created (where $j=1..k$).
FactReasoner 2 (FR2).
This version also retrieves up to $k$ contexts for each atom $A_{i}$, but it subsequently removes any duplicated contexts, thus resulting in $m$ unique contexts denoted by $\{C_{1},...,C_{m}\}$. It then considers the relationships between atom $A_{i}$ and all $m$ contexts, creating the factors $f(A_{i},C_{j})$, where $j=1..m$.
FactReasoner 3 (FR3).
In this version, we consider the same contexts $\{C_{1},...,C_{m}\}$ as in FR2, but in addition to the atom-context relationships we also consider the context-context relationships. Thus, we create the factors $f(A_{i},C_{j})$ as well as the factors $f(C_{j},C_{k})$, where $j=1..m$ and $k=1..m$ and $j\neq k$, respectively.
@@ 4 Experiments
Baseline Assessors.
Datasets.
We experimented with the following benchmark datasets: Biographies (Bio) Min et al. (2023), AskHistorians (AskH) Fangyuan Xu and Choi (2023), ELI5 Fangyuan Xu and Choi (2023), FreshBooks (Books) Song et al. (2024) and LongFact-Objects (LFObj) Wei et al. (2024).
@@
The Biographies is a labeled dataset that contains 157 biographies generated by ChatGPT for various person entities that have a Wikipedia page. Each biographic passage is also associated with a set of human generated atomic units (facts) that are labeled as supported (S) or not-supported (NS) by human annotators. We assume that this annotation is the ground truth.
@@
The other four datasets are unlabeled and consist of collections of prompts (or questions). Specifically, the AskH and ELI5 datasets contain 200 questions each that were scraped from the reddit/AskHistorians and reddit/explainlikeimfive online forums, while the Books dataset consists of 10 paragraphs sampled from 20 non-fictional books that were published between 2023 and 2024, for a total of 200 paragraphs. Our version of the LFObj dataset is a subset of the original dataset Wei et al. (2024) and contains 10 prompts sampled randomly from the original ones about objects spreading 38 different topics. For each prompt in these datasets, we generated a long-form response spanning up to two paragraphs using the llama-3.3-70b-instruct model Touvron et al. (2023).
Measures of Performance.
@@
For each dataset $\mathcal{D}$ and each competing assessor, we report the factual precision $Pr$ and $F_{1}@K$ measure, averaged over the number of prompts in $\mathcal{D}$. If $\mathcal{D}$ contains annotated atomic units (i.e., ground truth) then we also report the standard $F_{1}$ measure and the mean absolute error (MAE) given by:
@@
where $Pr_{j}$ and $Pr^{*}_{j}$ are the precision and respectively the true factual precision of the $j$-th instance. Since the FactReasoner assessors calculate the posterior marginal distributions of the atoms, we also compute the $\mathcal{E}$-measure. Finally, we also report the mean number of supported (#S), contradicted (#C) and undecided atoms (#U), respectively.
@@
External Knowledge Sources.
We consider two external knowledge sources: Wikipedia and Google Search results. For a given atom, the top $k$ results are retrieved as contexts either from wikipedia.org using the Wikipedia retriever available from LangChain [^6] https://python.langchain.com, or from google.com using the Serper API [^7] https://serper.dev. In both cases, a context is a tuple $(t,l,s,d)$, where $t$ is the title of the wiki/web-page, $l$ is the link, $s$ is a short text snippet or summary and $d$ is the content retrieved from $l$ (but capped at max 4000 characters). We used $k=3$ for the Wikipedia retriever and $k=5$ for the Google Search results Min et al. (2023); Wei et al. (2024).
In order to ensure a consistent evaluation, we decompose each response in the datasets into the corresponding atomic units (and subsequently revise them) using the same llama-3.3-70b-instruct model. Furthermore, we also retrieve and cache the relevant contexts for atoms from the two knowledge sources. This way, all competing assessors could be evaluated on the same sets of atoms and contexts. Table 2 summarizes the properties of the datasets, showing the number of prompts, the mean number of atoms and the median number of atoms ($K$). The latter is used for calculating the $F_{1}@K$ measure. In addition, for the labeled dataset, we also indicate the true number of supported atoms ($S^{*}$) and the true precision (Pr).
[…]
BERT-based relation model: albert-xlarge-vitaminc-mnli
FR1 12 5 12 0.40 0.66 0.39 0.25 0.11
FR2 10 16 4 0.32 0.53 0.31 0.34 0.09
FR3 10 16 4 0.32 0.53 0.31 0.33 0.09
LLM-based relation model: llama-3.1-70b-instruct
FR1 13 1 16 0.41 0.70 0.41 0.23 0.10
FR2 19 2 9 0.60 0.83 0.59 0.11 0.06
FR3 19 2 9 0.60 0.83 0.59 0.11 0.06
4.1 Evaluating the Relation Model
@@ 4.2 Results on Labeled Datasets
Table 4 shows the results obtained on the labeled Biographies dataset using Wikipedia retrieved contexts (the best performance is highlighted). We see that in terms or mean absolute error (MAE), precision and $F_{1}$ scores, the FR2 and FR3 assessors powered by stronger LLMs like llama-3.1-70b-instruct and mixtral-8x22b-instruct achieve the best performance compared to the other assessors. This is because both FR2 and FR3 can exploit the relationships between the atoms and all the retrieved contexts (as well as between the contexts themselves for FR3), not just the ones between an atom and its corresponding top $k$ contexts. Therefore, it is often the case that a context retrieved for atom $A_{i}$ may support or contradict another atom $A_{j}$ for which it wasn’t retrieved. This leads to a higher number of true positives and consequently larger $F_{1}$ scores. We also observe that the numbers of undecided atoms is also smaller for FR2/FR3 compared with the other assessors. The performance of FR3 is similar to that of FR2 because most of the context-context relationships are equivalence.
@@
We notice that the lightweight FR1 assessor performs on par with FV and VS in terms of precision, error and $F_{1}$ score. This shows that using only the top $k$ contexts to determine whether an atom is supported or not is fairly limited. Furthermore, in situations when an atom is supported by several contexts but is contradicted by another context (which might as well be a spurious contradiction), the FR2/FR3 assessors are able to correctly label the atom as supported based on the strengths of the respective relationships (i.e., probabilities of entailment and contradiction) whereas the other assessors struggle and often label the atom as contradicted or undecided. This demonstrates clearly the power of the probabilistic approach to factuality as employed by the proposed assessors.
Assessor # S # C # U Pr $\uparrow$ $F_{1}@K$uparrow $\mathcal{E}$downarrow
granite-3.0-8b-instruct
FS 17 5 0.76 0.74
FB 8 0 13 0.35 0.36
FV 12 4 5 0.55 0.55
FR1 (ours) 4 1 16 0.19 0.19 0.14
FR2 (ours) 10 9 2 0.46 0.47 0.09
FR3 (ours) 11 8 2 0.47 0.48 0.10
llama-3.1-70b-instruct
FS 15 7 0.69 0.68
FB 8 0 13 0.37 0.38
FV 5 0 16 0.25 0.25
FR1 (ours) 5 0 17 0.21 0.22 0.13
FR2 (ours) 10 1 10 0.45 0.46 0.09
FR3 (ours) 10 1 10 0.44 0.45 0.09
mixtral-8x22b-instruct
FS 16 6 0.71 0.70
FB 9 0 12 0.43 0.43
FV 7 0 14 0.34 0.34
FR1 (ours) 5 0 17 0.22 0.23 0.12
FR2 (ours) 11 0 11 0.46 0.47 0.09
FR3 (ours) 11 0 11 0.46 0.47 0.09
DeepSeek-v3 9 1 12 0.43 0.43
Table 5: Results on the unlabeled AskH dataset using Wikipedia contexts (mean number of supported (#S), contradicted (#C) and undecided (#U) atoms).
@@
Table 6: Results on the unlabeled AskH dataset using Google Search contexts (mean number of supported (#S), contradicted (#C) and undecided (#U) atoms).
@@ 4.3 Results on Unlabeled Datasets
Tables 5 and 6 show the results obtained on the unlabeled AskH dataset using Wikipedia and Google Search retrieved contexts, respectively (we include in the the Appendix the experiments with the remaining datasets: Books, ELI5 and LFObj). Since there is no ground truth for this dataset, we only report the precision, $F_{1}@K$ (for $K=22$) and the $\mathcal{E}$ measure. However, for reference, we also experimented with DeepSeek-v3, perhaps one of the strongest open models at the moment, using a suitable prompt DeepSeek-AI (2024).
We note that the AskH dataset covers a wider range of topics compared with the Biographies dataset and, therefore, the Wikipedia based contexts have a much smaller coverage in this case compared with the Google search results. This is reflected by the relatively smaller number of atoms supported by Wikipedia only contexts (Table 5) compared with those supported by Google Search results (Table 6), across all competing assessors. A similar pattern can be observed for the precision, $F_{1}@K$ and $\mathcal{E}$ measures reported in Table 5, as they are typically inferior to those shown in Table 6.
The prompt-based assessors FV and VS are fairly conservative in this case as well and find relatively fewer supported atoms compared with the FR2 and FR3 assessors. The latter two benefit from considering the relationships between an atom and all of the retrieved contexts and, therefore, find more supported atoms. The corresponding precision and $F_{1}@K$ values are also higher for FR2/FR3. We also observe that the $\mathcal{E}$-measure specific to the FR assessors correlates well with the number of supported atoms, namely as the number of supported atoms increases $\mathcal{E}$ gets closer to 00 .
When looking at the FS assessor, we notice again that it tends to find more supported atoms than the other assessors. However, we hypothesise that some of these atoms are false positive as before, but acknowledge that without any ground truth information it is difficult to verify this hypothesis.
Comparing the results obtained with DeepSeek-v3, we see that FV and VS come very close, especially for Google Search contexts. This is likely because the prompts used are fairly similar. In contrast, FR2/FR3 find slightly more supported atoms, although the results are very close for Wikipedia only contexts. We believe that this is caused by spurious context-atom entailment relationships which indicates that a better relation model is required.
In summary, our proposed FactReasoner assessor achieved the best performance on the labeled dataset, nearly matching the ground truth. However, on the unlabeled datasets, its performance was comparable with that of its competitors including DeepSeek-v3, a very powerful open model.
@@ 5 Related Work
The assessment of LLMs’ adherence to factual knowledge has gained significant attention in recent years due to their widespread adoption. Several well-established benchmarks, including TruthfulQA Lin et al. (2022), FreshQA Vu et al. (2023), HaluEval Li et al. (2023), HalluQA Cheng et al. (2023), and FELM Chen et al. (2023), focus on short-form response evaluation, where an LLM’s knowledge is tested through individual factoids classified as either true or false. More recent studies Min et al. (2023); Wei et al. (2024); Bayat et al. (2025); Song et al. (2024) have extended this approach to long-form generations by decomposing responses into distinct factual elements, which are then evaluated separately against relevant evidence retrieved from an external source of knowledge. These previous works typically assume that the retrieved pieces of information do not overlap or conflict with each other.
Conflicting information is prevalent in external knowledge sources Xu et al. (2024) and it typically impacts modern retrieval augmented-generation systems that aim to reduce hallucinations in LLMs Lewis et al. (2021). Other works have developed new benchmarks for capturing knowledge conflicts in realistic situations Hou et al. (2024); Marjanović et al. (2024); Su et al. (2024); Pham et al. (2024).
Our work is closely related with recent studies on self-consistency that aim at improving the logical consistency of the LLM’s response with respect to the input query by leveraging various methods including formal reasoning Wang et al. (2023); Dohan et al. (2022); Mitchell et al. (2022).

### v2

@@ 3.3 The FactReasoner Pipeline
The proposed FactReasoner pipeline for factuality assessment is shown in Figure 3 and consists of four main stages called Atomizer, Reviser, Retriever and Evaluator, respectively. It takes as input a response $y$ and outputs the marginal posterior probabilities $P(a_{i})$ of $y$’s atomic units together with the factuality measures described earlier, such as $Pr(y)$, $F_{1}@K(y)$ and $\mathcal{E}(y)$, respectively.
@@
FactReasoner 1 (FR1). In this case, for each atom variable $A_{i}$ up to $k$ most relevant contexts $\{C_{1}^{i},...,C_{k}^{i}\}$ are retrieved and only the relationships between each atom $A_{i}$ and its corresponding contexts are considered, namely only the factors $f(A_{i},C_{j}^{i})$ are created (where $j=1..k$).
FactReasoner 2 (FR2). This version also retrieves up to $k$ contexts for each atom $A_{i}$, but it subsequently removes any duplicated contexts, thus resulting in $m$ unique contexts denoted by $\{C_{1},...,C_{m}\}$. It then considers the relationships between each atom $A_{i}$ and all $m$ contexts, creating the factors $f(A_{i},C_{j})$, where $j=1..m$.
FactReasoner 3 (FR3). We consider the same contexts $\{C_{1},...,C_{m}\}$ as in FR2, but in addition to the atom-context relationships we also consider the context-context relationships. Thus, we create the factors $f(A_{i},C_{j})$ and $f(C_{j},C_{k})$, where $j=1..m$, $k=1..m$ and $j\neq k$, respectively.
@@ 4 Experiments
In this section, we empirically evaluate the FactReasoner assessor for long-form factuality and compare it against state-of-the-art approaches on labeled and unlabeled datasets. Although the FactReasoner pipeline stages can be instantiated with different LLMs, in our implementation we use the same LLM throughout the entire pipeline and focus our empirical evaluation on the Evaluator stage (i.e., factuality assessment).
Baseline Assessors. For our purpose, we consider the following state-of-the-art prompt-based long-form factuality assessors: FactScore (FS) Min et al. (2023), FactVerify (FV) Bayat et al. (2025) and VeriScore (VS) Song et al. (2024). FactScore is one of the first assessor that prompts an LLM to assess whether an atomic unit of the response is supported or not by a set of contexts relevant to the atom which are retrieved from an external knowledge source such as Wikipedia. FactVerify and VeriScore are more recent refinements of FactScore’s original prompt that can accommodate other external knowledge sources such as Google Search results and enable the LLM’s reasoning capabilities to evaluate the relationships between an atom and its relevant contexts. Unlike FactScore, the latter can label the atoms as supported, contradicted and undecided, respectively. In our experiments, we instantiated the competing assessors including the FactReasoner variants with open-source LLMs belonging to the IBM Granite [^3] https://huggingface.co/ibm-granite, Meta LLaMA [^4] https://huggingface.co/meta-llama and MistralAI Mixtral [^5] https://huggingface.co/mistralai families, namely: granite-3.0-8b-instruct, llama-3.1-70b-instruct, and mixtral-8x22b-instruct, respectively. All our LLMs are hosted remotely on compute nodes with A100 80GB GPUs and accessed via litellm APIs capable of serving 1500 prompts per second.
Datasets. We experimented with the following datasets: Biographies (Bio) Min et al. (2023), AskHistorians (AskH) Xu et al. (2023), ELI5 Xu et al. (2023), FreshBooks (Books) Song et al. (2024), and LongFact-Objects (LFObj) Wei et al. (2024). These datasets have been widely adopted in prior work and are considered representative benchmarks for long-form factuality assessment, as they encompass a diverse range of topics and tasks, including creative writing, history, astronomy, chemistry, and more.
@@
The Biographies is the only labeled dataset available. It contains 157 biographies generated by ChatGPT for various person entities that have a Wikipedia page. Each biographic passage is also associated with a set of human generated atomic units (facts) that were labeled as supported (S) or not-supported (NS) by human annotators. We assume that this annotation is the ground truth.
@@
The AskH, ELI5, Books and LFObj datasets are unlabeled and consist of collections of prompts. Specifically, the AskH and ELI5 datasets each contain 200 questions sourced from the Reddit forums r/AskHistorians and r/explainlikeimfive, respectively. The Books dataset comprises 200 paragraphs, sampled as 10 excerpts from each of 20 non-fiction books published between 2023 and 2024. Our version of the LFObj dataset is a curated subset of the original collection Wei et al. (2024), consisting of 10 prompts randomly selected from those related to objects spanning 38 distinct topics. For each prompt across these datasets, we generated a long-form response – up to two paragraphs in length – using the llama-3.3-70b-instruct model Touvron et al. (2023).
Additionally, we constructed a new dataset, Conflicts, comprising 1,000 claims (or atomic units) randomly sampled from the recent ConflictBank benchmark Su et al. (2024). Each claim, originally extracted from Wikidata, is considered true (i.e., supported). For every claim, we include two associated contexts: one supporting context (the default in ConflictBank) and one conflicting context (representing misinformation, also provided by ConflictBank). Notably, these two contexts are mutually contradictory, thus offering a controlled setting for our long-form factuality evaluation.
@@
Measures of Performance. For each dataset $\mathcal{D}$ and each competing assessor, we report the factual precision ($Pr$) and the $F_{1}@K$ score, averaged over all prompts in $\mathcal{D}$. If $\mathcal{D}$ includes annotated atomic units (i.e., ground truth labels), we additionally report the standard $F_{1}$ score and the mean absolute error (MAE), defined as:
@@
where $Pr_{j}$ and $Pr^{*}_{j}$ are the precision and the ground-truth precision for the $j$-th instance, respectively. Since the FactReasoner assessors calculate the posterior marginals of the atoms, we also compute the $\mathcal{E}$-measure. Finally, we include the mean number of atoms classified as supported (#S), contradicted (#C), and undecided (#U).
@@
External Knowledge Sources. We consider two external knowledge sources: Wikipedia and Google Search results. For a given atom, the top $k$ results are retrieved as contexts either from wikipedia.org using the Wikipedia retriever available from LangChain [^6] https://python.langchain.com, or from google.com using the Serper API [^7] https://serper.dev. In both cases, a context is a tuple $(t,l,s,d)$, where $t$ is the title of the wiki/web-page, $l$ is the link, $s$ is a short text snippet or summary and $d$ is the content retrieved from $l$ (but capped at max 4000 characters). We used $k=3$ for the Wikipedia retriever and $k=5$ for the Google Search results Min et al. (2023); Wei et al. (2024).
To ensure a consistent evaluation across all datasets, we decompose each generated response into its constituent atomic units and revise them using the same llama-3.3-70b-instruct model. Additionally, we retrieve and cache the relevant contextual information for each atom from the two designated knowledge sources. This standardized setup allows all competing assessors to be evaluated on an identical set of atoms and associated contexts. Table 2 summarizes the key properties of the datasets, including the number of prompts, the mean number of atoms per response, and the median number of atoms ($K$), which is used in computing the $F_{1}@K$ metric. For the labeled datasets, we also report the true number of supported atoms ($S^{*}$) and ground-truth precision (Pr).
@@ 4.1 Results on Labeled Datasets
Assessors Pr $F_{1}$ MAE Pr $F_{1}$ MAE
llama-3.1-70b-instruct mixtral-8x22b-instruct
FR2 vs FS 0.3916 0.0000 0.0001 0.0421 0.0000 0.0003
FR2 vs FV 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
FR2 vs VS 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Table 4: Statistical significance tests: $p$-values for Pr, $F_{1}$ and MAE obtained on the labeled Biographies dataset.
Biographies Dataset. Table 3 shows the results obtained on the labeled Biographies dataset using Wikipedia retrieved contexts (the best performance is highlighted). We see that in terms or mean absolute error (MAE), precision and $F_{1}$ scores, the FR2 and FR3 assessors powered by stronger LLMs like llama-3.1-70b-instruct and mixtral-8x22b-instruct achieve the best performance compared to the other assessors. This is because both FR2 and FR3 can exploit the relationships between the atoms and all the retrieved contexts (as well as between the contexts themselves for FR3), not just the ones between an atom and its corresponding top $k$ contexts. Therefore, it is often the case that a context retrieved for atom $A_{i}$ may support or contradict another atom $A_{j}$ for which it wasn’t retrieved. This leads to a higher number of true positives and consequently larger $F_{1}$ scores. We also observe that the numbers of undecided atoms is also smaller for FR2/FR3 compared with the other assessors. FR3 performs similarly to FR2 because the majority of the context-context relationships are equivalences.
@@
[…]
Additionally, we conducted one-sided $t$-tests on the $Pr$, $F_{1}$, and MAE metrics obtained by FR2 and its competitors, using the stronger LLaMA and Mixtral models. The resulting $p$-values are reported in Table 4. The near-zero $p$-values for the $F_{1}$ and MAE metrics indicate that FR2 significantly outperforms its competitors on these measures. In contrast, the relatively higher $p$-values observed between FR2 and FS suggest that FS exhibits higher precision but also a substantially higher false positive rate compared to FR2.
Assessor llama (70b) $\uparrow$ mixtral (22b) $\uparrow$ granite (8b) $\uparrow$
FS 0.35 0.74 0.49
FV 0.33 0.45 0.63
VS 0.06 0.46 0.56
FR1/2 (ours) 0.88 0.83 0.61
FR3 (ours) 0.83 0.89 0.62
Table 5: Accuracy on the labeled Conflicts dataset.
Conflicts Dataset. Table 5 shows the accuracy – defined as the proportion of claims correctly classified as true – obtained by the competing assessors on the Conflicts dataset. In this case, FR1 and FR2 are identical. Notably, the FR assessors consistently outperform the prompt-based methods, particularly when leveraging more powerful models such as LLaMA or Mixtral. For instance, FR2 and FR3 – both of which exploit the strength of supporting and conflicting relationships between the claims and their contexts – correctly classify over 80% of the claims when using the LLaMA model. In contrast, prompt-based approaches struggle when conflicts are present both between the contexts and between the claim and its contexts, resulting in significantly lower accuracy. These results underscore the effectiveness of the probabilistic reasoning framework employed by the proposed assessors in handling conflicts.
@@
Table 6: Results for the unlabeled AskH dataset using Google Search contexts (mean number of supported (#S), contradicted (#C) and undecided (#U) atoms).
Method # S # C # U Pr $\uparrow$ $F_{1}@K$uparrow $\mathcal{E}$downarrow
granite-3.0-8b-instruct
FS 20 2 0.87 0.84
FV 16 0 6 0.71 0.70
VS 18 2 3 0.76 0.75
FR1 (ours) 18 0 3 0.79 0.77 0.04
FR2 (ours) 21 1 0 0.90 0.86 0.02
FR3 (ours) 17 5 0 0.74 0.72 0.04
llama-3.1-70b-instruct
FS 20 3 0.84 0.82
FV 18 0 4 0.78 0.76
VS 17 0 5 0.72 0.71
FR1 (ours) 14 1 7 0.62 0.62 0.07
FR2 (ours) 19 1 3 0.80 0.78 0.04
FR3 (ours) 18 2 2 0.80 0.78 0.04
mixtral-8x22b-instruct
FS 20 3 0.84 0.82
FV 18 0 4 0.76 0.74
VS 18 0 4 0.79 0.77
FR1 (ours) 16 0 6 0.69 0.68 0.06
FR2 (ours) 20 0 2 0.86 0.83 0.03
FR3 (ours) 20 0 2 0.86 0.83 0.03
DeepSeek-v3 17 3 5 0.72 0.69
Table 7: Results for the unlabeled Books dataset using Google Search contexts (mean number of supported (#S), contradicted (#C) and undecided (#U) atoms).
@@ 4.2 Results on Unlabeled Datasets
Tables 6 and 7 show the results obtained on the unlabeled AskH and Books datasets using Google Search retrieved contexts, respectively (the Appendix includes the remaining datasets). Since there is no ground truth for these datasets, we only report the precision, $F_{1}@K$ (for $K=22$) and the $\mathcal{E}$-measure. However, for reference, we also experimented with DeepSeek-v3 DeepSeek-AI (2024), perhaps one of the strongest open models at the moment, using a suitable prompt (see Appendix).
The prompt-based assessors, FV and VS, are relatively conservative in this case and identify fewer supported atoms compared to the FR2 and FR3 assessors. In contrast, FR2 and FR3 benefit from evaluating the relationships between each atom and all retrieved contexts, enabling them to identify more supported atoms. This advantage is reflected in their higher precision and $F_{1}@K$ scores. We also observe that the $\mathcal{E}$-measure, specific to the FR assessors, correlates well with the number of supported atoms: as the number of supported atoms increases, $\mathcal{E}$ tends to approach to 0. Interestingly, the FS assessor, identifies more supported atoms than any other method. However, we hypothesize that a portion of these atoms may be false positive – an issue observed in the labeled datasets. Nonetheless, in the absence of ground-truth annotations, this hypothesis remains difficult to verify.
Compared to DeepSeek-v3, FV and VS yield very similar results – likely due to the similarity of their prompts. In contrast, FR2/FR3 identify slightly more supported atoms, though the difference is minimal. This is because some contexts support atoms they weren’t originally retrieved for.
In summary, our proposed FactReasoner assessor achieved the best performance on the labeled datasets, nearly matching the ground truth. However, on the unlabeled datasets, its performance was comparable with that of its competitors including DeepSeek-v3, a very powerful open model.
@@ 5 Related Work
Factuality evaluation of LLMs has received growing attention due to their widespread use. Early benchmarks such as TruthfulQA Lin et al. (2022), FreshQA Vu et al. (2023), HaluEval Li et al. (2023), HalluQA Cheng et al. (2023), and FELM Chen et al. (2023) focus on short-form factuality, assessing isolated factoids. More recent work Min et al. (2023); Wei et al. (2024); Bayat et al. (2025); Song et al. (2024) extends this to long-form responses by decomposing them into atomic facts evaluated against external evidence – typically assuming non-conflicting sources.
However, conflicting information is common in real-world knowledge bases Xu et al. (2024), posing challenges for retrieval-augmented generation systems Lewis et al. (2021). New benchmarks have emerged to capture such conflicts more realistically Hou et al. (2024); Marjanović et al. (2024); Su et al. (2024); Pham et al. (2024).
Our work is also related to recent efforts on improving self-consistency in LLMs through formal reasoning Wang et al. (2023); Dohan et al. (2022); Mitchell et al. (2022).